OpenStudy (anonymous):

Find the GCF. 18x3 and 30x5

jhonyy9 (jhonyy9):

do you know what mean GCF ?

jhonyy9 (jhonyy9):

great common factor ok ? so than what will be the GCF there ?

jhonyy9 (jhonyy9):

18x^3 = 3*6*x^3 30x^5 =5*6*x^3 *x^2 so not you check these and see what are common ?

jhonyy9 (jhonyy9):

6 is ?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

3 is ?

jhonyy9 (jhonyy9):

not

jhonyy9 (jhonyy9):

5 is ?

jhonyy9 (jhonyy9):

not

jhonyy9 (jhonyy9):

than the fisrt common factor will be 6

jhonyy9 (jhonyy9):

and continue

jhonyy9 (jhonyy9):

x^3 is common ?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

x^2 is common ?

jhonyy9 (jhonyy9):

not

jhonyy9 (jhonyy9):

than because there not are again terms so you can writing that the GCF = 6x^3 ok ?

jhonyy9 (jhonyy9):

hope that is understandably